# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ 10. Conclusion
Our work targets the automatic selection of pragma configurations for HLS with a framework that automatically inserts Merlin pragmas into loop-based programs. Our framework is guided by an analytical performance and resource model, which serves as a lower bound estimation for the achievable performance across all possible configurations. By formulating this model as a Non-Linear Program (NLP), the theoretically optimal pragma configuration can be determined. Our framework facilitates efficient design-space exploration by leveraging the latency lower bound property, allowing for the rapid elimination of points in the search space using a lightweight DSE process.
Appendix A Problem Size
{adjustwidth}
-1in-1in Kernel $\mathcal{O}(ops)$ $\mathcal{O}(Mem)$ Large Medium Small 2mm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 800, NJ 900, NK 1100, NL 1200 NI 180, NJ 190, NK 210, NL 220 NI 40, NJ 50, NK 70, NL 80 3mm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 800, NJ 900, NK 1000, NL 1100, NM 1200 NI 180, NJ 190, NK 200, NL 210, NM 220 NI 40, NJ 50, NK 60, NL 70, NM 80 adi $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 1000 TSTEPS 100, N 200 TSTEPS 40, N 60 atAx $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ M 1900, N 2100 M 390, N 410 M 116, N 124 bicg $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ M 1900, N 2100 M 390, N 410 M 116, N 124 cholesky $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 correlation $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1200, N 1400 M 240, N 260 M 80, N 100 covariance $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1200, N 1400 M 240, N 260 M 80, N 100 deriche $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ W 4096, H 2160 W 720, H 480 W 192, H 128 doitgen $\mathcal{O}(n^{4})$ $\mathcal{O}(n^{2})$ NQ 140, NR 150, NP 160 NQ 40, NR 50, NP 60 NQ 20, NR 25, NP 30 durbin $\mathcal{O}(n^{2})$ $\mathcal{O}(n)$ N 2000 N 400 N 120 floyd-warshall $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2800 N 500 N 180 ftdt-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TMAX 500, NX 1000 NY 1200 TMAX 100, NX 200, NY 240 TMAX 40, NX 60 , NY 80 gemm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 1000, NJ 1100, NK 1200 NI 200, NJ 220, NK 240 NI 60, NJ 70, NK 80 gemver $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 gesummv $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 1300 N 250 N 90 gramschmidt $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80 heat-3d $\mathcal{O}(n^{4})$ $\mathcal{O}(n^{3})$ TSTEPS 500, N 120 TSTEPS 100, N 40 TSTEPS 40, N 20 jacobi-1d $\mathcal{O}(n^{2})$ $\mathcal{O}(n)$ TSTEPS 500, N 2000 TSTEPS 100, N 400 TSTEPS 40, N 120 jacobi-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 1300 TSTEPS 100, N 250 TSTEPS 40, N 90 lu $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 ludcmp $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 mvt $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 nussinov $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2500 N 500 N 180 seidel-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 2000 TSTEPS 100, N 400 TSTEPS 40, N 120 symm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80 syr2k $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80 syrk $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80 trisolv $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120 trmm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80
@@ Appendix A Problem Size
Appendix B Theoretical Latency and Resource Modeling
We now outline key elements of our analytical performance model, and the associated proofs this model computes a lower bound on latency under resource constraints.
B.1. A Formal Model for Latency
Definition B.1 (Straight-line code).
Definition B.2 (Live-in set).
Definition B.3 (Live-out set).
Definition B.4 (Operation Graph).
Given a straight-line code region $R$ made of a list $L$ of statements $S\in L$, the operation graph OG is the directed graph <$<\{N,V_{I},root,V_{O}\},E>$ such that $\forall S_{i}\in L,~{}N_{S_{i}}\in N$; and for every operation with output $o$ and inputs $\vec{i}$ in $L$ $\forall S_{i},\forall i_{k}\in\vec{i}_{S_{i}},e_{i_{k},o}\in E$, $\forall S_{i}:(o_{S_{i}},{\vec{i}}_{S_{i}})\in L,S_{j}:(o_{S_{j}},{\vec{i}}_{S_{j}})\in L$ with $S_{i}\neq S_{j}$ then we have $E_{S_{i},S_{j}}\in E$ iff $o_{Si}\cap{\vec{i}}_{S_{j}}\neq\emptyset$. For every input (resp. output) in $S_{i}$ which is not matched with an output (resp. input) of another $S_{j}$ in $L$, create a node $V_{val}\in V_{I}$ (resp. $V_{O}$) for this input (resp. output) value. If $dim({\vec{i}}_{S_{i}})=0$ then an edge $e_{root,S_{i}}$ is added to $E$.
From this representation, we can easily define key properties to subsequently estimate the latency and area of this code region, such as its span, or critical path.
Definition B.5 (Operation Graph critical path).
B.1.1. Latency Lower Bound
We can build a simple a lower bound on the latency of an operation graph:
Theorem B.6 (Lower bound on latency of an Operation Graph).
Proof Th B.6.
Every operation $S_{i}$ is associated with at least one edge with a source in <$<\{V_{I},root\}>$, so there is a path between one of these nodes and every operation by construction in Def. B.4. For an operation to produce a useful output, there must be a path from its output to a node in $V_{O}$, otherwise the operation may be removed by dead code elimination. Therefore the shortest path $sp$ between a pair of nodes $(v_{i},v_{o})\in<\{V_{I},root\},V_{O}>$ is the shortest sequence of operations in dependence that must be executed to produce $v_{o}$. As any operation takes at least one cycle to complete per Def. B.6, then this path must take at least #$\#sp$ cycles to complete. As we take the largest of the shortest paths between all possible pairs $(v_{i},v_{o})$ then $OG_{cp}^{L}$ is the length of the longest shortest path to produce any output $v_{o}$ from some input, via a sequence of producer-consumer operations. Therefore it must take at least #$\#OG_{cp}^{L}$ cycles to execute this path. ∎
Theorem B.7 (Latency Lower Bound under Operation Resource Constraints).
Given $R_{op}$ a count of available resources of type $op$, for each operation type, and $LO(op)$ the latency function for operation $op$, with $LO(op)\geq 1$. #$\#L(op)$ denotes the number of operations of type $op$ in $L$. We define $LO(\#OG_{cp}^{L})=\sum_{n\in cp}LO(n)$ the critical path weighted by latency of its operations. The minimal latency of a region $L$ is bounded by
$Lat_{R_{op}}^{L}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil\#L(o)\times LO(o)/R_{o}\rceil))$
Proof Th B.7.
Suppose $\forall o\in op,~{}R_{o}\geq\#L(o)$. Then there is equal or more resources available than work to execute, this is equivalent to the infinite resource hypothesis of Th. B.6, the minimal latency is $LO(\#OG_{cp}^{L})$.
This theorem provides the building block to our analysis: if reasoning on a straight-line code region, without any loop, then building the operation graph for this region and reasoning on its critical path is sufficient to provide a latency lower bound.
B.1.2. Loop Unrolling: full unroll
It follows a simple corollary:
Corollary B.8 (Equivalence between fully unrolled and straight-line code).
Given a loop nest $l$, if full unrolling is applied to $l$ then the code obtained after full unrolling is a straight-line code as per Def. B.1.
Proof Co B.8.
By construction the process of fully unrolling all the loops creates a straight-line code region without loop control, which therefore fits Def. B.1. ∎
Consequently, we can bound the latency of a fully unrolled loop nest:
Theorem B.9 (Minimal latency of a fully unrolled loop nest).
$Lat_{R_{op}}^{l}\geq\max(LO(\#OG_{cp}^{L}),\max_{o\in op}(\lceil\#L(o)\times L_{o}/R_{o}\rceil))$
Proof Th B.9.
By Corollary B.8. ∎
B.1.3. Loop Unrolling: partial unroll
Unrolling can be viewed as a two-step transformation: first strip-mine the loop by the unroll factor, then consider the inner loop obtained to be fully-unrolled. The latency of the resulting sub-program is determined by how the outer-loop generated will be implemented. We assume without additional explicit information this unrolled loop will execute in a non-pipelined, non-parallel fashion. Note this bound requires to build the operation graph for the whole loop body. This is straightforward for inner loops and/or fully unrolled loop nests, but impractical if the loop body contains other loops. We therefore define a weaker, but more practical, bound that enables composition:
Theorem B.10 (Minimal latency of a partially unrolled loop with factor UF).
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor\times Lat_{R_{op}}^{L^{{}^{\prime}}}$
Proof Th B.10.
By construction and Theorem B.7, $Lat_{R_{op}}^{L^{{}^{\prime}}}$ is a lower bound on the latency of $L^{{}^{\prime}}$, that is the sub-program made of $UF$ iterations of the loop. $\lfloor TC/UF\rfloor\leq TC_{l}/UF$ is a lower bound on the number of iterations of the loop. As we assume a non-pipelined execution for the resulting outer loop, every iteration shall start after the completion of the preceding one, that is its iteration latency, itself bounded by $Lat_{R_{op}}^{L^{{}^{\prime}}}$. ∎
Theorem B.11 (Minimal latency of a partially unrolled loop with factor UF and complex loop bodies).
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor*Lat_{R_{op}}^{L}$
Proof Th B.11.
Vitis allows to do a reduction with a tree reduction in logarithmic time with the option “unsafe-math”.
Theorem B.12 (Minimal latency of a partially unrolled loop with factor UF for reduction loop with tree reduction).
Proof Th B.12.
By definition a reduction loop is a loop with a dependency distance of 1. Hence, at each iteration the same memory cell is read and write. Because of the dependency distance of 1, only one element can be added to the same memory cell. However each data can be adding independently two by two and the result of this independent addition can also be adding two by two until we obtained one value. Hence the reduction can be done in $\log_{2}(UF)$ iterations with a tree reduction. As the depth of the tree is $\log_{2}(UF)$ and each node at the same depth can be executed independently in $Lat_{R_{op}}^{L}$ cycles, the straight line code has a latency greater or equal to $Lat_{R_{op}}^{L}\times\lfloor\log_{2}(UF)\rfloor$. And then similarly to Th. B.11 and B.12 the sequential execution of the loops without pragma repeat this process $\lfloor TC/UF\rfloor$ times.
1L1: for (i = 0; i < 8; i++) {
2 c += a[i];
3}
Listing 12: Example of code demonstrating a reduction, where a tree reduction technique can be applied, as depicted in Figure 1.
B.1.4. Loop pipelining
Loop pipelining amounts to overlapping multiple iterations of the loop, so that the next iteration can start prior to the completion of the preceding one. The initiation interval (II) measures in cycles the delay between the start of two consecutive iterations. It is easy to prove our formula template accurately integrates the latency of pipelined loops with the $I$ operator. We compute the minimal II in function of the dependencies of the pipelined loop and the iteration latency of the operations of the statements during the NLP generation. Let $RecMII$ and $ResMII$ be the recurrence constraints and the resource constraints of the pipelined loop, respectively. We have $II\geq\max{(ResMII,RecMII)}$. $RecMII=\max_{i}\lceil\frac{delay(c_{i})}{distance(c_{i})}\rceil$ with $delay(c_{i})$ the total latency in dependency cycle $c_{i}$ and $distance(c_{i})$ the total distance in dependency cycle $c_{i}$. We suppose that $ResMII=1$, as we do not know how the resource will be used by the compiler. Hence, if the loop is a reduction loop then the $II\geq\frac{IL_{reduction}}{1}$ with $IL_{reduction}$ the iteration latency of the operation of reduction.
It follows a bound on the minimal latency of a pipelined loop:
Theorem B.13 (Minimal latency of a pipelined loop with known II).
$Lat_{R_{op}}^{l,P}\geq Lat_{R_{op}}^{L}+II*(TC_{l}-1)$
Proof Th B.13.
$Lat_{R_{op}}^{L}$ is the minimal latency to complete one iteration of $l$ by Theorem B.7. The initiation interval measures the number of elapsed cycles before the next iteration can start, it takes therefore at least $TC_{l}*II-1$ to start $TC_{l}-1$ iterations, irrespective of their completion time. Therefore the latency of the loop is at least the latency of one iteration to complete, and for all iterations to be started. ∎
B.1.5. Loop pipelining and unrolling
Theorem B.14 (Minimal latency of a pipelined loop with known II and partially unrolled).
$Lat_{R_{op}}^{l,P}\geq Lat_{R_{op}}^{L^{{}^{\prime}}}+II*(\frac{TC_{l}}{UF}-1)$
Proof Th B.14.
By construction and Theorem B.9 the latency $Lat_{R_{op}}^{L^{{}^{\prime}}}$ is a lower bound of $L^{{}^{\prime}}$. As the loop was split due to the partial unrolling, the trip count of the pipelined loop is $\frac{TC_{l}}{UF}$. Theorem B.13 gives us the lower bound of the latency for a loop with a trip count equal to $\frac{TC_{l}}{UF}$.
B.1.6. Non-Parallel, Non-Pipelined Loops
Definition B.15 (Lower bound on latency of a non-parallel, non-pipelined loop under resources constraints).
$Lat_{R_{op}}^{l}\geq TC_{l}*Lat_{R_{op}}^{L}$
B.1.7. Coarse-Grained parallelization
Theorem B.16 (Minimal latency of coarse-grained unrolled loop).
$Lat_{R_{op}}^{l,S}\geq\lfloor TC/UF\rfloor\times Lat_{R_{op}}^{L^{{}^{\prime}}}$
Proof Th B.16.
By construction, Definition B.15, Theorem B.13 and definition of the loop body $L$, $Lat^{L}_{R_{op}}$ is a lower bound of the loop body $L$. As the loop is not a reduction loop there is no dependency between the loop bodies of $L$ for different iteration of $l$ and then the loop bodies can be executed in parallel. If $UF<TC_{l}$, then $\lfloor TC_{l}/UF\rfloor\leq TC_{l}/UF$ is a lower bound on the number of iterations of the loop. As we assume a non-pipelined execution for the resulting outer loop, every iteration shall start after the completion of the preceding one, that is its iteration latency, itself bounded by $Lat_{R_{op}}^{L^{{}^{\prime}}}$.
B.1.8. Program latency lower bound under resource constraints
Theorem B.17.
Proof Th B.17.
Given a program and the available resource of DSP $DSP_{avail}$, if $R^{min}_{used}<DSP_{avail}$ the lower bound is valid and the program does not over-utilize the resources.
B.1.9. Memory transfer
Theorem B.18 (Lower bound of the memory transfer latency for an array).
Given a loop body $L$, the set of array $\mathcal{A}$, an array $a\in\mathcal{A}$, and $LAT^{mem}_{a}$ the latency to transfer the array $a$ from off-chip to on-chip (inputs) and from on-chip to off-chip (outputs). $\forall a\in\mathcal{A},LAT^{mem}_{a}\geq(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$. With $\mathds{1}_{cond}=1$ if $cond$ = true else 0.
Proof Th B.18.
In order to transfer all the elements of the array $a$ we can use packing with a maximum packing allowed by the target device of $max\_burst\_size$, which means in practice the real burst size will be equal or less than $max\_burst\_size$. As all the elements of the array $a$ are in the same bank, the transfer is sequential. And as we as suppose all operation including memory transfers are done in at least one cycle, the minimum latency is $footprint_{a}/max\_burst\_size$ to transfer once the array $a$. As an array can be input, output or both we need to add the transfer from off-chip to on-chip for inputs i.e., $a\in V_{I}^{L}$ and from on-chip to off-chip for the outputs i.e., $a\in V_{O}^{L}$.
Theorem B.19 (Lower bound of the memory transfer latency).
Given a loop body $L$, the set of array $\mathcal{A}$, the number of cycles to transfer the array $a$ is bounded by $\max_{a\in\mathcal{A}}(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$.
Proof Th B.19.
According to Th. B.18 the lower bound to transfer one array $a$ is $(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$. As the array can be on different DRAM banks the transfer from off-chip to on-chip can be done in parallel but at least one array has a latency greater or equal to $(\mathds{1}_{a\in V_{O}^{L}}+\mathbb{\in V_{I}^{L}})\times footprint_{a}/max\_burst\_size$ and hence the memory latency is equal to $\max_{a\in\mathcal{A}}(\mathds{1}_{a\in V_{O}^{L}}+\mathds{1}_{a\in V_{I}^{L}})\times footprint_{a}/burst\_size$. ∎
B.2. Summary
By composing all the theorems, this allows us to end up with the final latency lower bound of the program which is presented in theorems B.20 for the computation and B.21 for the computation and communication.
Theorem B.20 (Computation Latency Lower bound of a Program).
with
Proof Th B.20.
Through composition and the application of Theorems B.9, B.10 and B.11, $Lat^{L_{fg}}_{R_{op}}$ serves as a computation latency lower bound for the fully unrolled sub-loop body of $L$, denoted as $L_{fg}$, where $Lat_{L_{par}}+Lat_{L_{seq}}\times\prod_{l\in\mathcal{L}^{reduction}_{L_{fg}}}\frac{TC_{l}}{UF_{l}}\times\log_{2}(UF_{l})$ represent the critical path of $Lat^{L_{fg}}_{R_{op}}$.
By employing composition alongside Theorems B.13 and B.14, $Lat^{L_{pip}}_{R_{op}}$ stands as a computation latency lower bound for $L_{pip}$.
Utilizing composition, Theorem B.16, and Definition B.15, $Lat^{L}_{R_{op}}$ emerges as a computation latency lower bound for $L$.
Theorem B.21 (Latency lower bound of a program optimized with Merlin pragmas).
Given available resource $DSP_{avail}$ and a program which contains a loop body $L$ with a computation latency $Lat_{L}^{computation}$ and a communication latency $Lat_{L}^{communication}$.
$Lat_{L}=Lat_{L}^{computation}+Lat_{L}^{communication}$
Proof Th B.21.
The AMD/Xilinx Merlin compiler does not overlap computation and communication. Hence the computation and communication is a sum of the latency of computation and latency of communication. By composition and Theorems B.20 and B.19.

### v4

@@ 11. Conclusion
In this paper, we presented NLP-DSE, a framework that automates the selection of pragma configurations for high-level synthesis (HLS) by inserting Merlin pragmas into loop-based programs. The core of our approach is an analytical performance and resource model, which serves as a lower bound estimation for achievable performance across all pragma configurations. By leveraging this model and formulating it as a Non-Linear Program (NLP), our framework can identify the theoretically optimal pragma configurations.
Our method significantly improves design-space exploration (DSE) efficiency by using the latency lower bound property to quickly eliminate suboptimal design points. This enables a lightweight DSE process that drastically reduces exploration time without sacrificing the quality of results (QoR). Compared to existing approaches like AutoDSE and HARP, NLP-DSE achieves equal or better QoR with fewer design points explored and in significantly less time.
Extensive evaluations on a variety of benchmarks have shown that NLP-DSE consistently delivers superior or equivalent results while maintaining fast exploration times.
Now that our framework can automatically insert pragma configurations, the next step is to extend its capabilities to include code transformations. Transformations like loop permutation, tiling, and data reuse strategies can further optimize performance by restructuring the code to better exploit hardware resources. Integrating these transformations with pragma insertion will allow our framework to explore a broader range of design optimizations, potentially unlocking even greater improvements in both performance and resource utilization within the HLS process.
Acknowledgments
This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.
Appendix A Proofs for Latency and Resource Lower Bound
A.1. Latency Lower Bound
Proof Th 6.6.
Every operation $S_{i}$ is associated with at least one edge with a source in <$<\{V_{I},root\}>$, so there is a path between one of these nodes and every operation by construction in Def. 6.4. For an operation to produce a useful output, there must be a path from its output to a node in $V_{O}$, otherwise the operation may be removed by dead code elimination. Therefore the shortest path $sp$ between a pair of nodes $(v_{i},v_{o})\in<\{V_{I},root\},V_{O}>$ is the shortest sequence of operations in dependence that must be executed to produce $v_{o}$. As any operation takes at least one cycle to complete per Def. 6.6, then this path must take at least #$\#sp$ cycles to complete. As we take the largest of the shortest paths between all possible pairs $(v_{i},v_{o})$ then $OG_{cp}^{L}$ is the length of the longest shortest path to produce any output $v_{o}$ from some input, via a sequence of producer-consumer operations. Therefore it must take at least #$\#OG_{cp}^{L}$ cycles to execute this path. ∎
Proof Th 6.7.
Suppose $\forall o\in op,~{}R_{o}\geq\#L(o)$. Then there is equal or more resources available than work to execute, this is equivalent to the infinite resource hypothesis of Th. 6.6, the minimal latency is $LO(\#OG_{cp}^{L})$.
A.1.1. Loop Unrolling: full unroll
Proof Co 6.8.
By construction the process of fully unrolling all the loops creates a straight-line code region without loop control, which therefore fits Def. 6.1. ∎
Proof Th 6.9.
By Corollary 6.8. ∎
A.1.2. Loop Unrolling: partial unroll
Proof Th 6.10.
By construction and Theorem 6.7, $Lat_{R_{op}}^{L^{{}^{\prime}}}$ is a lower bound on the latency of $L^{{}^{\prime}}$, that is the sub-program made of $UF$ iterations of the loop. $\lfloor TC/UF\rfloor\leq TC_{l}/UF$ is a lower bound on the number of iterations of the loop. As we assume a non-pipelined execution for the resulting outer loop, every iteration shall start after the completion of the preceding one, that is its iteration latency, itself bounded by $Lat_{R_{op}}^{L^{{}^{\prime}}}$. ∎
Proof Th 6.11.
Proof Th 6.12.
By definition a reduction loop is a loop with a dependency distance of 1. Hence, at each iteration the same memory cell is read and write. Because of the dependency distance of 1, only one element can be added to the same memory cell. However each data can be adding independently two by two and the result of this independent addition can also be adding two by two until we obtained one value. Hence the reduction can be done in $\log_{2}(UF)$ iterations with a tree reduction. As the depth of the tree is $\log_{2}(UF)$ and each node at the same depth can be executed independently in $Lat_{R_{op}}^{L}$ cycles, the straight line code has a latency greater or equal to $Lat_{R_{op}}^{L}\times\lfloor\log_{2}(UF)\rfloor$. And then similarly to Th. 6.11 and 6.12 the sequential execution of the loops without pragma repeat this process $\lfloor TC/UF\rfloor$ times.
A.1.3. Loop pipelining
Proof Th 6.13.
$Lat_{R_{op}}^{L}$ is the minimal latency to complete one iteration of $l$ by Theorem 6.7. The initiation interval measures the number of elapsed cycles before the next iteration can start, it takes therefore at least $TC_{l}*II-1$ to start $TC_{l}-1$ iterations, irrespective of their completion time. Therefore the latency of the loop is at least the latency of one iteration to complete, and for all iterations to be started. ∎
A.1.4. Loop pipelining and unrolling
Proof Th 6.14.
By construction and Theorem 6.9 the latency $Lat_{R_{op}}^{L^{{}^{\prime}}}$ is a lower bound of $L^{{}^{\prime}}$. As the loop was split due to the partial unrolling, the trip count of the pipelined loop is $\frac{TC_{l}}{UF}$. Theorem 6.13 gives us the lower bound of the latency for a loop with a trip count equal to $\frac{TC_{l}}{UF}$.
A.1.5. Coarse-Grained parallelization
Proof Th 6.16.
By construction, Definition 6.15, Theorem 6.13 and definition of the loop body $L$, $Lat^{L}_{R_{op}}$ is a lower bound of the loop body $L$. As the loop is not a reduction loop there is no dependency between the loop bodies of $L$ for different iteration of $l$ and then the loop bodies can be executed in parallel. If $UF<TC_{l}$, then $\lfloor TC_{l}/UF\rfloor\leq TC_{l}/UF$ is a lower bound on the number of iterations of the loop. As we assume a non-pipelined execution for the resulting outer loop, every iteration shall start after the completion of the preceding one, that is its iteration latency, itself bounded by $Lat_{R_{op}}^{L^{{}^{\prime}}}$.
A.1.6. Program latency lower bound under resource constraints
Proof Th 6.17.
A.1.7. Memory transfer
Proof Th 6.18.
In order to transfer all the elements of the array $a$ we can use packing with a maximum packing allowed by the target device of $max\_burst\_size$, which means in practice the real burst size will be equal or less than $max\_burst\_size$. As all the elements of the array $a$ are in the same bank, the transfer is sequential. And as we as suppose all operation including memory transfers are done in at least one cycle, the minimum latency is $\frac{footprint_{a}}{max\_burst\_size}$ to transfer once the array $a$. As an array can be input, output or both we need to add the transfer from off-chip to on-chip for inputs i.e., $a\in V_{I}^{L}$ and from on-chip to off-chip for the outputs i.e., $a\in V_{O}^{L}$.
Proof Th 6.19.
According to Theorem 6.18, the lower bound for transferring one array $a$ is given by $\frac{footprint_{a}}{max\_burst\_size}$. If the array $a$ is transferred within a loop $l$, we only need to transfer the footprint of the array $a$ accessed within that loop multiplied by the trip count of the loops that iterate over this array and over the loop $l$, which corresponds to $footprint_{a}$ in the best case because we need to transfer at least all the elements of $a$ one time. As the array can reside on different DRAM banks, the transfer from off-chip to on-chip can be performed in parallel if transferred consecutively, i.e., at the same level. However, at least one array will have a latency greater than or equal to $\frac{footprint_{a}}{max\_burst\_size}$ so the latency of communication under the loop $l$ is bounded by $\max_{a\in\mathcal{A}}(loop_{l}\_cache\_array\_a\times\frac{footprint_{a}}{max\_burst\_size})$. Since communication cannot be overlapped if the transfers are not consecutive (i.e., not under the same loops), we sum the latency of communication for each loop. ∎
A.2. Summary
Proof Th 6.20.
Through composition and the application of Theorems 6.9, 6.10 and 6.11, $Lat^{L_{fg}}_{R_{op}}$ serves as a computation latency lower bound for the fully unrolled sub-loop body of $L$, denoted as $L_{fg}$, where $Lat_{L_{par}}+Lat_{L_{seq}}\times\prod_{l\in\mathcal{L}^{reduction}_{L_{fg}}}\frac{TC_{l}}{UF_{l}}\times\log_{2}(UF_{l})$ represent the critical path of $Lat^{L_{fg}}_{R_{op}}$. By employing composition alongside Theorems 6.13 and 6.14, $Lat^{L_{pip}}_{R_{op}}$ stands as a computation latency lower bound for $L_{pip}$. Utilizing composition, Theorem 6.16, and Definition 6.15, $Lat^{L}_{R_{op}}$ emerges as a computation latency lower bound for $L$.
Proof Th 6.21.
The AMD/Xilinx Merlin compiler executes computation and communication sequentially, making total latency their sum. Applying composition and Theorems 6.20 and 6.19 derives a lower bound on latency.
Appendix B Evaluation vs. HARP
@@ A.2. Summary
Appendix C Problem Size
Kernel $\mathcal{O}(ops)$ $\mathcal{O}(Mem)$ Large Medium Small
2mm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 800, NJ 900, NK 1100, NL 1200 NI 180, NJ 190, NK 210, NL 220 NI 40, NJ 50, NK 70, NL 80
3mm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 800, NJ 900, NK 1000, NL 1100, NM 1200 NI 180, NJ 190, NK 200, NL 210, NM 220 NI 40, NJ 50, NK 60, NL 70, NM 80
adi $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 1000 TSTEPS 100, N 200 TSTEPS 40, N 60
atAx $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ M 1900, N 2100 M 390, N 410 M 116, N 124
bicg $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ M 1900, N 2100 M 390, N 410 M 116, N 124
cholesky $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
correlation $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1200, N 1400 M 240, N 260 M 80, N 100
covariance $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1200, N 1400 M 240, N 260 M 80, N 100
deriche $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ W 4096, H 2160 W 720, H 480 W 192, H 128
doitgen $\mathcal{O}(n^{4})$ $\mathcal{O}(n^{2})$ NQ 140, NR 150, NP 160 NQ 40, NR 50, NP 60 NQ 20, NR 25, NP 30
durbin $\mathcal{O}(n^{2})$ $\mathcal{O}(n)$ N 2000 N 400 N 120
floyd-warshall $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2800 N 500 N 180
ftdt-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TMAX 500, NX 1000 NY 1200 TMAX 100, NX 200, NY 240 TMAX 40, NX 60 , NY 80
gemm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ NI 1000, NJ 1100, NK 1200 NI 200, NJ 220, NK 240 NI 60, NJ 70, NK 80
gemver $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
gesummv $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 1300 N 250 N 90
gramschmidt $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80
heat-3d $\mathcal{O}(n^{4})$ $\mathcal{O}(n^{3})$ TSTEPS 500, N 120 TSTEPS 100, N 40 TSTEPS 40, N 20
jacobi-1d $\mathcal{O}(n^{2})$ $\mathcal{O}(n)$ TSTEPS 500, N 2000 TSTEPS 100, N 400 TSTEPS 40, N 120
jacobi-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 1300 TSTEPS 100, N 250 TSTEPS 40, N 90
lu $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
ludcmp $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
mvt $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
nussinov $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ N 2500 N 500 N 180
seidel-2d $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ TSTEPS 500, N 2000 TSTEPS 100, N 400 TSTEPS 40, N 120
symm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80
syr2k $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80
syrk $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80
trisolv $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$ N 2000 N 400 N 120
trmm $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$ M 1000, N 1200 M 200, N 240 M 60, N 80